OpenStudy (anonymous):

write an equation for the sequence 6,7.2,8.64

OpenStudy (aum):

7.2 / 6 = 1.2 8.62 / 7.2 = 1.2 This is a geometric sequence with common ratio 1.2. \(\large a_n = 6 * (1.2)^{n-1}\)

OpenStudy (anonymous):

thank you

OpenStudy (aum):

yw.

OpenStudy (shinalcantara):

i wonder if this is in geometric progression since if we try to really have the exact quotient, 8.62/7.2 is 1.197222... for you to have an exact quotient of 1.2, the third term must be 8.64.. hmmm

OpenStudy (anonymous):

oops my bad it is 8.64